OpenStudy (anonymous):

Find the point on the terminal side of θ = negative three pi divided by four that has an x coordinate of -1. Show your work for full credit.

OpenStudy (anonymous):

I am really confused, this lesson wasn't descriptive and I don't even know where to start.

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@triciaal

OpenStudy (triciaal):

terminal side means where it ends draw a grid and mark where x = -1 angles are measured from the right horizontal in an anti-clockwise direction can you sketch this using the draw tool so I can see it? x = r cos theta = -1 (y = r sin theta)

OpenStudy (triciaal):

a complete circle = 2 Pi

OpenStudy (anonymous):



OpenStudy (anonymous):

I know this is not complete Im just confuse. Im sorry.

OpenStudy (triciaal):

that was correct now mark the angle

OpenStudy (anonymous):

What I dont understand is how you got the y point.

OpenStudy (michele_laino):

hint: here is the angle \(\theta= - 3\pi/4\)